Please help me solve this question. What are the dimensions of a parallelogram if the base is 10 more than twice the height and the area is 208 cm�? Thank you.
Answer by CubeyThePenguin(3113)

area of parallelogram = (base)(height)
height = h

(h)(2h + 10) = 208
2h^2 + 10h - 208 = 0
h^2 + 5h - 104 = 0
(h - 8)(h + 13) = 0

h = 8, h = -13

Distance can't be negative, so the height is 8 cm and the base is 26 cm.